OpenStudy (anonymous):

Angle A is a supplement of the angle B and angle C is a complement of the angle B. What is the difference between angles A and C?

OpenStudy (amistre64):

supple = 180 comps = 90

OpenStudy (amistre64):

a+b = 180 b+c = 90 b = 90-c ; and b = 180-a 90-c = 180-a a-c = 180-90 = 90

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

(: